# PSA-MIL: A Probabilistic Spatial Attention-Based Multiple Instance Learning for Whole Slide Image Classification

### v1

@@ 3.4 Diversity Loss for Spatial Decay Parameters
To encourage variation among heads, we introduce an entropy-based loss term that promotes diversity in $\{\theta_{h}\}_{h=1}^{H}$. The entropy of $p(\theta)$ is given by: $H(p)=-\mathbb{E}_{\theta\sim p(\theta)}[\log p(\theta)]$, which we estimate via Monte Carlo [16]:
@@
Figure 4: Performance comparison of different approaches for incorporating spatial context into self-attention MIL (MSA-MIL).
Figure 5: AUC on TCGA-CRC vs. FLOPs (log scale) for contextual MIL models, with bubble size representing parameter count.
@@ 4 Experiments
Datasets We validated our findings on two cancer subtyping tasks: Microsatellite Stable (MSS) vs. Microsatellite Instability (MSI) for colorectal adenocarcinoma (CRC) [11, 1] and Genomic Stability (GS) vs. Chromosomal Instability (CIN) for stomach adenocarcinoma (STAD) [39, 44]. Both datasets were sourced from The Cancer Genome Atlas [^1] https://portal.gdc.cancer.gov (TCGA) project. For details on clinical significance, slide counts, and label distribution, refer to Peled et al. [34].
In line with previous studies [13, 26, 42, 36], we assessed model performance primarily using the Area Under the Curve (AUC) metric, with Accuracy and F1 as secondary measures. To ensure robustness, we obtained results using nested 5-fold cross-validation, repeated three times with different seeds.
Implementation Details All experiments used the foundation backbone from [19]. PSA-MIL architecture comprises a single layer with three 32-dim heads, $\tau=10^{-3}$, and learning rate $10^{-4}$. Experiments were conducted on an NVIDIA A100 GPU. Please refer to the supplementary for more details.
Figure 6: (A) Tumor segmentation using NCT-CRC-HE-100K [21]. (B-D) Attention heatmaps of individual heads in PSA-MIL[Gau], illustrating diverse spatial focus. (E) Aggregated attention heatmap from the attention pooling operator.
Figure 7: Ablation studies on the impact of $\alpha$ (a-b) and $\tau$ (c-d) in PSA-MIL.
4.1 Benchmarking with Existing Works
We compare PSA-MIL to several contextual and non-contextual MIL baselines, concentrating on recent advancements in spatial modeling, such as SM-MIL [4], Bayes-MIL [9], and GTP [46]. The results are presented in Tab. 1.
@@ 4.1 Benchmarking with Existing Works
Our PSA-MIL variants consistently outperform prior works across most metrics, demonstrating the effectiveness of learned spatial priors. Notably, non-contextual methods perform competitively compared to previous contextual approaches, with IBMIL even surpassing certain contextual models. This suggests that prior contextual methods may not fully capture spatial relationships effectively. Among contextual methods, SM-MIL demonstrates strong performance, highlighting the effectiveness of local smoothness in attention maps. Similarly, Bayes-MIL performs well, particularly on TCGA-STAD, underscoring the advantages of probabilistic modeling. Furthermore, we observe performance differences among PSA-MIL variants, suggesting that spatial distribution modeling impacts performance.
4.2 Evaluating Spatial Context Integration in Self-Attention MIL
@@
To further demonstrate the effectiveness of our approach, we compare it to several existing methods for integrating spatial context into self-attention MIL (MSA-MIL), described in Sec. 3.1.1. Specifically, we benchmark against three baselines: the naive approach (with no spatial context), K-local attention (where only instances within the K-neighborhood are included in the self-attention, each with equal contribution), PPEG used in TransMIL [36], and Relative Positional Encoding (RPE) suggest by SETMIL [45].
@@
Fig. 4 presents the results. PSA-MIL achieved the best results on both datasets, outperforming the baseline by +5.5% in TCGA-CRC and +3.1% in TCGA-STAD. While positional encoding methods provided some improvement, their impact was significantly lower than ours, with RPE proving more effective than PPEG. The K-local attention variants yielded minimal to no performance gains, highlighting the limitations of fixed local constraints and the advantages of dynamic spatial modeling.
4.3 Computational Cost Analysis of Contextual MIL Models
@@
Contextual MIL models inherently incur higher computational costs due to the need to model spatial interactions. To further demonstrate the superiority of our approach, we conduct a computational cost analysis, comparing AUC performance against FLOPs, with bubble sizes indicating the number of trained parameters. For this experiment, we utilized TCGA-CRC dataset. Fig. 5 presents the results.
@@
PSA-MIL achieves the highest AUC while maintaining the lowest FLOP count and fewer parameters, indicating the effectiveness of our proposed spatial pruning strategy. SM-MIL showed good performance (Tab. 1) but requires significantly more computational resources. GTP has a small parameter footprint yet suffers from high FLOPs and inferior performance.
@@ 4.4 Ablation Studies
4.4.1 Impact of $\alpha$ Value
Our approach employs diversity loss to encourage the capture of diverse spatial patterns. As shown in Fig. 7 (a-b), TCGA-CRC favored relatively higher $\alpha$ values, indicating the benefits of diverse representations. On the other hand, a moderate $\alpha$ value yields the best results across all decay functions for TCGA-STAD.
4.4.2 Impact of $\tau$ Value
Our approach leverages $\tau$ to regulate spatial pruning, controlling the range of spatial interactions in PSA-MIL. In our experiments, we typically set it to a constant value (e.g., $1e{-3}$), which we find works well in practice. As shown in Fig. 7 (c-d), Gaussian decay, which models gradual changes, remains relatively stable to variations in $\tau$. In contrast, Cauchy decay, which favors long-range dependencies, is more sensitive, exhibiting steeper performance drops. Despite the expected performance drops, PSA-MIL maintains strong results even under significant computational constraints.
4.5 Training Dynamics and Attention Visualization
We analyze the training dynamics and attention patterns of PSA-MIL[Gau] on the TCGA-CRC dataset, focusing on the impact of diversity loss and the learned spatial behaviors of attention heads.
Fig. 3 examines the evolution of the local-K in our dynamic local attention formulation. Without diversity loss (Fig. 3a), all heads converge to similar locality values, whereas with diversity loss (Fig. 3b), they exhibit more distinct behaviors, ensuring diverse spatial representations. Similarly, inter-head similarity remains relatively high without diversity loss (Fig. 3c), while diversity loss (Fig. 3d) encourages head specialization by reducing pairwise similarity.
In Fig. 6, we visualize the attention heatmaps of PSA-MIL[Gau]. All attention maps align with tumor regions, demonstrating the model’s ability to identify relevant areas. Additionally, we observe distinct attention behaviors across heads: head 2 (K=4) exhibits highly localized attention with dispersed activation spots, whereas head 3 (K=17) captures global dependencies with a smoother attention distribution. This reinforces the role of learned spatial priors in producing complementary attention patterns.
5 Conclusion
We introduce PSA-MIL, a novel approach for integrating spatial context into attention-based MIL through a probabilistic interpretation of self-attention as a posterior distribution. Unlike conventional MIL methods, which often fail to fully exploit spatial dependencies, PSA-MIL dynamically learns spatial relationships during training. Additionally, our entropy-based diversity loss promotes distinct spatial attention patterns across heads, further enhancing representation learning. Our experiments show that PSA-MIL outperforms existing methods while significantly reducing computational costs, highlighting its effectiveness in adaptive spatial modeling.

### v2

@@ 3.3 Diversity Loss for Spatial Decay Parameters
To encourage variation among heads, we introduce an entropy-based loss term that promotes diversity in $\{\theta_{h}\}_{h=1}^{H}$. The entropy of $p(\theta)$ is given by: $H(p)=-\mathbb{E}_{\theta\sim p(\theta)}[\log p(\theta)]$, which we estimate via Monte Carlo:
@@
3.4 Spatially Pruning the Posterior
The quadratic complexity of self-attention poses a challenge for long sequences, such as in WSI analysis. Solutions like low-rank and kernel-based approximations often deliver suboptimal performance [12, 15].
To address this challenge, we propose a spatial pruning strategy for the posterior distribution that restricts attention computations based on learned spatial dependencies. By leveraging our prior distribution modeling, as described in Sec. 3.2.2, we are able to identify insignificant interactions before full computation, thereby achieving substantial cost reduction without compromising performance. Namely, we apply a filtering criterion $f(d|\theta)\geq\tau$, where $\tau$ is a threshold parameter, permitting only instances with decays above $\tau$ to contribute to the posterior computation. Such pruning techniques have been explored in mixture models, demonstrating the effectiveness of removing components that contribute negligibly to the posterior [37, 30].
By leveraging an invertible parametric function $f(d|\theta)$, we derive a guaranteed upper bound on spatial interactions: $d\leq f^{-1}(\tau|\theta)$. As a result, the computational complexity is reduced to $O(n\cdot K^{2})\ll O(n^{2})$, where $K=\lceil f^{-1}(\tau|\theta)\rceil$. This yields a dynamic local attention, where $K$ is learned via $\theta$, and the retained instances are weighted by their learned decay values, unlike fixed local attention. Fig. 3 (a-b) shows the evolution of our dynamic local attention during training. To the best of our knowledge, this is the first instance where an attention locality is derived during training in an end-to-end fashion. Notably, we do not fine-tune $\tau$; we typically set it to a constant value (e.g., 1e31e{-3}), which we find works well in practice. Fig. 2 illustrates how $f^{-1}(\tau|\theta)$ is applied in the PSA-MIL pipeline.
@@ 4 Experiments
We conducted extensive experiments spanning cancer subtyping, metastatic detection, survival prediction, and patch-level localization, with further analyses of computational efficiency, training dynamics, and attention visualizations. Unless otherwise specified, the Lunit foundation model [20] was used as our pretrained encoder. An NVIDIA A100 GPU was used for training. All evaluations were performed under identical conditions, including the same encoder, data splits, and official codebases. Additional details and extended results are provided in the supplementary material.
4.1 Cancer Subtyping
@@
We consider two subtyping tasks [32]: MSS vs. MSI in Colorectal Cancer (CRC) and CIN vs. GS in Stomach Cancer (STAD), using datasets from TCGA [^1] https://portal.gdc.cancer.gov. We compared PSA-MIL to several contextual and non-contextual MIL baselines, concentrating on recent advancements in spatial modeling, such as SM-MIL [5], Bayes-MIL [11]. Tab. 1 presents the results. Our PSA-MIL variants consistently outperform prior works across most metrics, demonstrating the effectiveness of learned spatial priors. Notably, non-contextual methods perform competitively compared to previous contextual approaches, with IBMIL even surpassing certain contextual models. This suggests prior contextual methods may capture inaccurate spatial relationships, leading to suboptimal performance. Among contextual methods, SM-MIL and Bayes-MIL performed well, highlighting the benefits of local smoothness and probabilistic modeling, respectively. Furthermore, we observe performance differences among PSA-MIL variants, highlighting the impact of spatial distribution modeling on performance.
4.1.1 Spatial Context Integration in Self-Attention MIL
@@
To further demonstrate the effectiveness of our approach, we compare it to several existing methods for integrating spatial context into self-attention MIL (MSA-MIL), described in Sec. 3.1.1. Specifically, we benchmark against five baselines: the naive approach (with no spatial context), K-local attention (where only instances within the K-neighborhood are included in the self-attention, each with equal contribution), PPEG used in TransMIL [35], Relative Positional Encoding (RPE) suggested by SETMIL [44], and RoPe 2D adaptation [36, 33].
@@
Tab. 2 presents the results. PSA-MIL achieved the best results on both datasets, surpassing the baseline by +5.5% in TCGA-CRC and +3.1% in TCGA-STAD. RoPE showed competitive gains on TCGA-CRC but performed poorly on TCGA-STAD. Prior work has noted that RoPE may underperform when sequence lengths vary [25, 9], and its 2D adaptation to pathology likely amplifies this limitation. Other positional encoding methods provided some improvement, but their impact was significantly lower than ours, with RPE proving more effective than PPEG. The K-local attention variants yielded minimal to no performance gains, highlighting the limitations of fixed local constraints and the advantages of dynamic spatial modeling.
MSA-MIL (baseline) Local [K=7] Local [K=15] PPEG RPE RoPe PSA-MIL [Gau]
TCGA-CRC 84.34 83.98 84.89 86.00 86.68 87.38* 88.97*
$\triangle$ $0.4\%$ $0.7\% +$$2.0\% +$$2.8\% +$$3.6\% +$$5.5\%$
TCGA-STAD 80.52 80.80 81.19 81.25 81.91 81.06 82.98
$\triangle$ $0.3\% +$$0.8\% +$$0.9\% +$$1.7\% +$$0.7\% +$$3.1\%$
Table 2: AUC scores and improvements over the MSA-MIL baseline, comparing different spatial modeling strategies. Significant gains over baseline (paired t-test, $p<0.05$) are marked with *.
4.2 Metastatic Detection
We evaluate our method on cancer metastatic detection using the CAMELYON16 dataset [1], assessing both slide-level classification and patch-level localization performance. Patch-level scores are extracted and compared to ground-truth tissue annotations only at test time. See the supplementary materials for more details. Results in Tab. 3 show that PSA-MIL achieved top slide-level classification performance, with AUC exceeding 96% and the highest F1 score. Crucially, it also excelled in localization, achieving the highest AUC-FROC and AUC-ROC by a clear margin. PSA-MIL stands out by maintaining high performance across both tasks, highlighting its enhanced robustness compared to other models.
Slide-level Patch-level
Method AUC- ROC F1 AUC- FROC AUC- ROC F1
ABMIL [19] 95.3 91.5 67.8 88.8 66.6
DTFD-MIL [42] 96.1 92.2 68.3 88.1 65.9
TransMIL [35] 94.8 90.5 67.7 84.6 21.1
GTP [45] 91.7 90.1 69.9 79.6 46.1
SM-MIL [5] 96.8 89.6 69.9 91.6 64.6
BAYES-MIL [11] 95.0 91.8 72.5 93.6 76.5
PSA-MIL[Gau] 96.1 92.3 75.9 94.7 75.4
Table 3: Slide-level classification and patch-level localization performance on CAMELYON16. PSA-MIL achieved top slide-level performance while maintaining superior localization quality.
Figure 4: Visual comparisons of different ROIs produced by various MIL models for metastatic detection. PSA-MIL accurately highlights the ROI; other methods show imprecise localization.
4.2.1 Visual Comparisons
The quantitative improvements in localization are further illustrated in Fig. 4, which shows heatmaps generated by different MIL baselines. PSA-MIL accurately spots the metastatic region, presenting a visually compelling and focused attention map. BAYES-MIL exhibited a more dichotomous attention distribution, with some regions of the metastatic area overlooked. SM-MIL, on the other hand, highlighted the entire ROI but produced a noisier attention map overall, likely due to over-smoothing. These findings emphasize PSA-MIL’s ability to generate precise and interpretable attention maps.
4.3 Survival Prediction
We evaluated our approach for survival prediction using TCGA-CRC, TCGA-STAD, and TCGA-BRCA cohorts, assessing performance with the concordance index (c-Index). We employed the survival cross-entropy loss as proposed in [41]. Results using UNI encoder [8] are presented in Tab. 4, with additional results using Lunit [20] and GigaPath [40] encoders provided in the supplementary materials.
PSA-MIL[Gau] achieved significantly better results on TCGA-CRC and TCGA-STAD, as well as the second-highest performance on TCGA-BRCA. Other contextual methods, such as Sm-MIL and BayesMIL, failed to achieve competitive scores. Surprisingly, ABMIL demonstrated enhanced performance, outperforming several of the contextual methods. These results remained consistent across encoders, highlighting the robustness of PSA-MIL.
Dataset AB- DTFD- Trans- Sm- Bayes- PSA-
MIL MIL MIL MIL MIL MIL[Gau]
TCGA-CRC 63.8 55.5 59.9 63.9 55.9 70.7
TCGA-STAD 58.6 53.3 52.9 56.6 52.4 61.1
TCGA-BRCA 61.9 53.9 57.8 57.2 49.9 60.2
Table 4: C-index results for survival prediction using the UNI encoder [8]. Results with Lunit [20] and GigaPath [40] (see supplementary) show consistent trends, confirming PSA-MIL’s superiority across encoders.
4.4 Analysis
4.4.1 Computational Cost Analysis
@@
Contextual MIL models inherently incur higher computational costs due to the need to model spatial interactions. To further demonstrate the superiority of our approach, we conduct a computational cost analysis in Fig. 5, comparing AUC performance against FLOPs, with bubble sizes indicating the number of trained parameters. We report mean FLOPs per batch during training, providing a consistent measure of computational cost across models.
@@
PSA-MIL achieves the highest AUC while maintaining the lowest FLOP count and fewer parameters, indicating the effectiveness of our proposed spatial pruning strategy. SM-MIL showed good performance (Tab. 1) but requires an order of magnitude more computational resources. GTP has a small parameter footprint yet suffers from high FLOPs and inferior performance.
4.4.2 Analyzing Training Dynamics
We analyze the training dynamics of PSA-MIL[Gau] on TCGA-CRC subtyping, focusing on the impact of diversity loss and the learned spatial attention patterns. Fig. 3 (a-b) examines the evolution of our dynamic local attention formulation. Without diversity loss, all heads converge to similar locality values, whereas with diversity loss, they exhibit more distinct behaviors, ensuring diverse spatial representations. Similarly, inter-head similarity remains relatively high without diversity loss (Fig. 3c), while diversity loss (Fig. 3d) encourages head specialization by reducing pairwise similarity. Additional visualizations of the attention heatmaps of PSA-MIL[Gau] are provided in the supplementary materials. These illustrate how different attention heads learn distinct spatial patterns. For example, the head with K=4K{=}4 produces highly localized activation spots, while the head with K=17K{=}17 captures broader, smoother attention distributions.
Figure 5: Subtyping performance (AUC) vs. mean FLOPs per batch during training (log scale) for contextual MIL models on TCGA-CRC, with bubble size representing parameter count. PSA-MIL delivered the best performance with a notably smaller computational footprint.
Figure 6: Ablation studies on the impact of diversity loss ($\alpha$) and pruning threshold ($\tau$) in PSA-MIL on TCGA-CRC. Extended results are provided in the supplementary material.
@@ 4.5 Ablation Studies
We report ablations on diversity loss weight ($\alpha$) and pruning threshold ($\tau$) in Fig. 6. Ablations indicate that higher $\alpha$ values generally improve performance, highlighting the beneficial effect of diverse representations. Results remain relatively stable across $\tau$, especially with Gaussian decay, enabling notable efficiency gains with minimal losses. Full analysis with additional results is provided in the supplementary.
5 Discussion
PSA-MIL introduces a probabilistic interpretation of self-attention that incorporates spatial context through learnable distance-decayed priors, an entropy-based diversity loss, and a spatial pruning strategy. This formulation enables attention heads to independently capture diverse locality ranges, improving both interpretability and computational efficiency. While effective, PSA-MIL is not without limitations. First, it remains bound to pretrained encoders, which can act as information bottlenecks if not chosen appropriately. Second, cross-domain evaluation, while highly valuable, is beyond the scope of this work and represents an important direction for future research. In summary, PSA-MIL combines strong performance with efficiency, paving the way for broader applicability.
6 Acknowledgments
M.F. is supported in part by research grants from the Israeli Ministry of Science and Technology (1001577479, 1001577564), the Israel Innovation Authority (80783), and the Zimin Institute for AI Solutions in Healthcare.
Y.E.M. is supported in part by the Israeli Science Foundation personal grant 2794/21 and the Israeli Personalized Medicine Program grant 3120/22.